OpenStudy (anonymous):

Write the word or phrase that best completes each statement or answers the question. Use Newton's method to find the positive fourth root of 3 by solving the equation x4 - 3 = 0. Start with x1 = 1 and find x2. please show me how to solve because its a short answer question. please don't make it confusing so i can rewrite. i cant afford to get this questions wrong

OpenStudy (anonymous):

its x^4-3=0

OpenStudy (anonymous):

\(f(1)=1-3=-2\) and \(f'(x)=4x^3\) so \(f'(1)=4\) your first guess is \(1\) your next guess is \[1-\frac{f(1)}{f'(1)}=1-\frac{-2}{4}=\frac{3}{2}\]

OpenStudy (anonymous):

so what does that mean?

OpenStudy (anonymous):

do i write all of that? and then the next step would be to solve an equation using 3/2?

OpenStudy (anonymous):

your first guess is \(x_1\) and your next guess is \[x_2=x_1-\frac{f(x_1)}{f'(x_1)}\]

OpenStudy (anonymous):

you were only asked for the next approximation, which in this case was \(\frac{3}{2}\)

OpenStudy (anonymous):

so that equation that equals 3/2 is the answer?

OpenStudy (anonymous):

so if i were to write this on a test everything you wrote at first would be correct? are u sure?

OpenStudy (anonymous):

so where does the fourth root of 3 come in?

OpenStudy (anonymous):

pretty sure, since you are only asked for the next approximation the one after that would be \[x_3=\frac{3}{2}-\frac{f(\frac{3}{2})}{f'(\frac{3}{2})}\] whatever that is

OpenStudy (anonymous):

keep going and you will get closer and closer

OpenStudy (anonymous):

so that is not the answer?

OpenStudy (anonymous):

the first answer i wrote is a complete answer to the question asked

OpenStudy (anonymous):

so why did u say keep going